Find the volume of the tetrahedron having the given vertices. (5, -6, -4), (-6, 4, -5), (4, -6, -3), (0, 0, 10)
Answer #1

A = (-5-6-4) B = (-6, 4,-5) C=(4,-6,-3) D= (0, 0, 10). A B = (-6, 4,-5) - (5, -6, 4) = (-11, 10, -1) Be = (4, -6, -3) - (-6,